# Spin Seebeck Effect as a Probe for Majorana Fermions in Kitaev Spin Liquids

### v1

@@ III.1 Real-time dynamics simulations
The contrasting behaviors between the KSL and FFM states can be ascribed to the fact that the carriers of the spin current are different between the two states. In the high-field FFM state, the elementary excitations are conventional magnons, which give a positive spin current irrespective of the details of the magnetic coupling. In contrast, in the low-field KSL state, the spin degree of freedom is fractionalized into itinerant Majorana fermions and localized $Z_{2}$ fluxes. They are independent of each other under magnetic fields, but in a low-field and low-temperature regime, the spin current in the KSL is expected to be carried predominantly by the itinerant Majorana fermions. The remarkable sign difference of $\tilde{J}_{S}$ between the FM and AFM Kitaev models suggests that the nature of Majorana fermions is different between the FM and AFM KSLs. This is a qualitatively different result from the one-dimensional antiferromagnet, where the spinon-mediated spin current was found to be negative [59]. We will discuss this sign switching by using the Majorana representation and the perturbation theory with respect to the magnetic field in Sec. III.2.
Let us make two remarks. The first one is on the magnitude of the tunnel spin current, $|\tilde{J}_{S}|$. In the KSL, $|\tilde{J}_{S}|$ increases much faster with temperature in the case of AFM than FM. Since the field strength $B$ is different and the low-field phase diagrams look different between the two cases, it is difficult to conclude that this is a general tendency. Nevertheless, $|\tilde{J}_{S}|$ are sufficiently large to be observed in experiments in both FM and AFM cases. This is explicitly shown by the comparison with the prototypical FM Heisenberg model, whose tunnel spin current is plotted by the black dotted line in Fig. 4(c) (see Appendix A for the details). Note that the spinon-mediated spin current was observed in the previous experiment [58], even though it is supposed to be smaller than the magnon-mediated one by a factor of $10^{-3}$$10^{-4}$. Meanwhile, in the FFM state, $|\tilde{J}_{S}|$ is almost the same at low temperature regime in FM and AFM. This is due to almost the same magnitude of Zeeman gap in FM and AFM, as shown in Figs. 5(a) and 5(b), which is further confirmed by the linear spin-wave theory in Sec. III.3. Figure 5(c) also shows $\tilde{J}_{S}$ for the FM Heisenberg model at $B=0.5$ for comparison; the magnitude is much smaller than the result for $B=0.01$ shown in Fig. 4(c) because of the larger Zeeman gap, but it is comparable to those for the FFM states in the FM and AFM Kitaev models.
@@ III.2 Perturbation analysis
To examine the origin of the different sign of $\tilde{J}_{S}$ between the FM and AFM KSLs in Fig. 4(c), we calculate the tunnel spin current in terms of Majorana fermions by using the perturbation theory with respect to the magnetic field. By using a Majorana fermion representation based on the Jordan-Wigner transformation [89, 90, 91], the Kitaev Hamiltonian at zero field in the flux free sector (the eigenstates without $Z_{2}$ flux excitations) is represented as
@@
with the Majorana fermion operators $c_{j}$. Based on the general framework of the perturbation theory [92], we obtain the effective operator of $S^{+}_{j_{0}}$ in the flux free sector through a transformation by $\Gamma:=PP_{0}(P_{0}PP_{0})^{-1/2}$, where $P_{0}$ and $P$ are projectors to the flux free sector of the unperturbed Hamiltonian [Eq. (1) for $B=0$] and to the states of the full Hamiltonian that are adiabatically connected to those in the flux free sector at $B=0$, respectively. By expanding up to the second order with respect to $B=|{\bf B}|$, the effective operator is obtained as
@@
$\left.X_{+}(\omega)\right|_{S^{\pm}_{j_{0}}\to s^{\pm}_{j_{0}},\mathcal{H}\to\mathcal{H}_{0}}=\delta\sum_{\ell}\frac{|\langle\ell|s^{-}_{j_{0}}|\emptyset\rangle|^{2}-|\langle\ell|s^{+}_{j_{0}}|\emptyset\rangle|^{2}}{(E_{\ell}-E_{\emptyset}-\omega)^{2}+\delta^{2}},$ (13)
@@
$\tilde{J}_{S}\sim\frac{B^{3}}{|K|^{3}}\tilde{j}_{S},$ (14)
@@
Figure 7: Results of the perturbation analysis for the FM and AFM Kitaev models in the weak magnetic field limit with ${\bf B}\parallel a$, ${\bf B}\parallel b$, and ${\bf B}\parallel c$: (a) $\omega$ dependences of $x_{+}(\omega)$ [Eq. (16)] and (b) temperature dependences of $\tilde{j}_{S}$ [Eq. (15)]. The lines represent the full results including all the contributions from $s^{+(1)}$ and $s_{1,2}^{+(2)}$, and the hatched regions represent the results ignoring $s^{\pm(2)}_{2}$. In the latter, ${\bf B}\parallel a$ and ${\bf B}\parallel b$ give the same results.
@@
The perturbation analysis provides a useful insight on the origin of the sign difference of the tunnel spin current between the FM and AFM cases. When we neglect the four Majorana contributions of $s_{2}^{\pm(2)}$, one can show that $x_{+}(\omega)$ in the FM and AFM cases are identical in the absolute value, but the signs are opposite. This is because the $\mathcal{O}(B^{3})$ contributions in Eq. (13) from the cross terms between $s^{\pm(1)}$ and $s^{\mp(2)}_{1}$ change their sign by $K\to-K$, i.e., $\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle\to-\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle$, due to the sign change in the eigenstates of Majorana fermions (see Appendix B). Consequently, the interchange of FM and AFM $K$ causes a sign reversal of $\tilde{j}_{S}$ in Eq. (15). The contributions with neglecting $s_{2}^{\pm(2)}$ are plotted by the hatched regions in Fig. 7. The result indicates that the contributions from the quadratic terms of the Majorana operators, $s^{\pm(1)}$ and $s^{\pm(2)}_{1}$, are dominant in $\tilde{j}_{S}$, and moreover, are the main cause of the sign switching of the tunnel spin current between the FM and AFM cases. The asymmetry between the two cases is due to the quartic terms in $s_{2}^{\pm(2)}$, leading to the faster growth of the tunnel spin current with temperature in AFM than FM, which is also consistent with the TDVP results in Fig. 4(c).
Let us make a remark on the $B$ dependence of the tunnel spin current. The perturbation theory in the low-field limit predicts $\tilde{j}_{S}\propto B^{3}$, but $\tilde{J}_{S}$ obtained by TDVP appears not to follow this asymptotic form; we find that $\tilde{J}_{S}$ at $B=B_{\rm KSL}/2$ is about half of that at $B=B_{\rm KSL}$ (not shown), suggesting that $\tilde{J}_{S}$ is roughly proportional to $B$. There are at least two possible reasons for this discrepancy. One is the contributions from excited states beyond the flux-free sector, which are neglected in the current perturbation analysis. The intermediate states [$|\ell\rangle$ in Eq. (13)] including flux excitations may give $\mathcal{O}(B)$ contributions, although they will be suppressed exponentially in the low-field limit because of the finite flux gap. If this is the case, the perturbation theory predict $\tilde{j}_{S}\simeq pB^{3}+qBe^{-\Delta_{\rm f}/T}$, where $\Delta_{\rm f}$ is the flux gap $\simeq 0.06|K|$ [24]. The other possibility is the limited system sizes in the TDVP simulations. Larger size simulations allow us to distinguish the $B$ dependence by accessing lower-field regions with higher precision by taking smaller dumping factor $\delta$ in Eq. (3). While such calculations require huge computational costs and are left for future study, this study comprehensively suggests that the coefficient $p$ and $q$ are both positive (negative) for the FM (AFM) KSL.
@@ III.3 Linear spin-wave theory
Figure 8: Results by the linear spin-wave theory for the (a,b) FM and (c,d) AFM Kitaev models at $B=B_{\rm FFM}$ with four field directions ${\bf B}\parallel a$, ${\bf B}\parallel b$, ${\bf B}\parallel c$, and ${\bf B}\parallel z$. (a) and (c) show the $\omega$ dependences of $X_{+}(\omega)$ for the FM and AFM cases, respectively. (b) and (d) show the temperature dependences of the tunnel spin current $\tilde{J}_{S}$.
@@
Figure 8 shows the results for $B=B_{\rm FFM}$ (see Fig. 3) with four different field directions. We find that $X_{+}(\omega)$ is always positive for both FM and AFM cases, as shown in Figs. 8(a) and 8(c), respectively. These behaviors are in qualitative agreement with the TDVP simulation results shown in Figs. 5(a) and 5(b). The positive $X_{+}(\omega)$ leads to the tunnel spin current with positive sign as shown in Fig. 8(b). This is also in agreement with the TDVP simulation results in Fig. 5(c). We note, however, that the gap in the spectrum of $X_{+}(\omega)$ is independent of the field directions in the linear spin-wave theory, while it is considerably different between ${\bf B}\parallel z$ and ${\bf B}\parallel c$ in the TDVP simulation.
@@ IV.1 Sign of spin current in Kitaev spin liquids
The real-time dynamics simulations based on TDVP reveal the sign switching of the tunnel spin current between the FM and AFM KSLs (Sec. III.1). This behavior is successfully reproduced by the perturbation theory in the Majorana representation (Sec. III.2). These findings indicate that the spin current in the KSL is predominantly carried by the itinerant Majorana fermions. More importantly, they suggest that the itinerant Majorana fermions contribute to the spin current with (up-)down-spin like nature in the (AFM) FM KSL where the sign of the tunnel spin current is (negative) positive, even though they do not have spin angular momentum.
The negative spin Seebeck effect that we find for the AFM KSL appears to be common to that for a quantum spin liquid in a quasi-one-dimensional antiferromagnet [58]. This implies common physics in the spin current carried by fractional excitations in AFM quantum spin liquids in one and two dimensions, even though the fractional quasiparticles have different nature: The spin angular momentum carried by a spinon in the one-dimensional quantum spin liquid is half quantized $\Delta S=1/2$, while that by a Majorana fermion in the KSL is not. Our results, however, indicate that the FM quantum spin liquid, which is peculiar to the Kitaev model, can induce a tunnel spin current with positive sign, suggesting that the spin Seebeck effect is useful to identify the nature of fractional quasiparticles in quantum spin liquids. We emphasize that our study reveals, for the first time to our knowledge, that the spin Seebeck effect in quantum spin liquids can be either positive or negative by analyzing both FM and AFM Kitaev models in two dimensions. Toward deeper understanding of the spin Seebeck effect in quantum spin liquids, further comparative studies are necessary for different types of quantum spin liquids in different dimensions. The simulation method for the real-time dynamics used in this study provides a powerful method to study generic models and awaits wider applications in the future.
Meanwhile, the spin Seebeck effect for the FM KSL has the same positive sign as that carried by conventional magnons in the FFM state. Thus, a way to distinguish between the two is needed. We will discuss this point in the next subsection.
@@ IV.2 Magnetic field dependence of tunnel spin current
Figure 9: Schematics of the magnetic field dependence of the tunnel spin current caused by the spin Seebeck effect at low temperature for (a) the FM and AFM Kitaev models and (b) the FM Heisenberg model.
@@
Based on the results by the real-time dynamics simulations, the perturbation theory, and the linear spin-wave theory, we show a schematic of the magnetic field dependence of the tunnel spin current caused by the spin Seebeck effect for the Kitaev model in Fig. 9(a). On one hand, in the low-field KSL region, the tunnel spin current increases in amplitude with increasing the field strength, with a different sign between the FM and AFM cases. As discussed in the end of Sec. III.2, a more detailed analysis is needed to fully elucidate the behavior in the low-field limit. On the other hand, in the high-field FFM regime, the tunnel spin current is positive for both FM and AFM cases and decreases with increasing the field strength as shown in Fig. 9(a). This is similar to the FM Heisenberg model [Fig. 9(b)]. Thus, these results imply that the tunnel spin current is maximized between the KSL and FFM states in the FM case, while it should change its sign in the AFM cases, as schematically shown in Fig. 9(a). However, further studies are needed to clarify the field dependence in those transient regions, including the another intermediate spin liquid region in the AFM case (see Fig. 3).
@@
Figure 10: Field angle dependences of the tunnel spin current when the magnetic field is rotated in the $ac$ plane obtained by (a) the perturbation analysis for low $B$ and (b) the linear spin-wave theory at $B=B_{\rm FFM}$. The angle $\theta$ is measured from the $a$ axis, and the data are calculated at $T=0.05$. In (a), the lines represent the full results including all the contributions from $s^{+(1)}$ and $s_{1,2}^{+(2)}$, and the hatched regions represent the results ignoring $s^{\pm(2)}_{2}$.

### v2

@@ III.1 Real-time dynamics simulations
The contrasting behaviors between the KSL and FFM states can be ascribed to the fact that the carriers of the spin current are different between the two states. In the high-field FFM state, the elementary excitations are conventional magnons, which give a positive spin current irrespective of the details of the magnetic coupling. In contrast, in the low-field KSL state, the spin degree of freedom is fractionalized into itinerant Majorana fermions and localized $Z_{2}$ fluxes. They are no longer independent of each other under magnetic fields, but in a low-field and low-temperature regime, the spin current in the KSL is expected to be carried predominantly by the itinerant Majorana fermions. The remarkable sign difference of $\tilde{J}_{S}$ between the FM and AFM Kitaev models suggests that the nature of Majorana fermions is different between the FM and AFM KSLs. This is a qualitatively different result from the one-dimensional antiferromagnet, where the spinon-mediated spin current was found to be negative [63]. We will discuss this sign switching by using the Majorana representation and the perturbation theory with respect to the magnetic field in Sec. III.2.
Let us make two remarks. The first one is on the magnitude of the tunnel spin current, $|\tilde{J}_{S}|$. In the KSL, $|\tilde{J}_{S}|$ increases much faster with temperature in the case of AFM than FM. Since the field strength $B$ is different and the low-field phase diagrams look different between the two cases, it is difficult to conclude that this faster increase of $|\tilde{J}_{S}|$ is a general tendency. Nevertheless, $|\tilde{J}_{S}|$ are sufficiently large to be observed in experiments in both FM and AFM cases. This is explicitly shown by the comparison with the prototypical $S=1/2$ FM Heisenberg model, whose tunnel spin current is plotted by the black dotted line in Fig. 4(c) (see Appendix A for the details). Note that the spinon-mediated spin current was observed in the previous experiment [62], even though it is supposed to be smaller than the magnon-mediated one by a factor of $10^{-4}$$10^{-3}$. We will further discuss in Sec. IV.4 the experimental observability by quantitative comparison with a typical three-dimensional (3D) ferromagnet in a realistic experimental setup. Meanwhile, in the FFM state, $|\tilde{J}_{S}|$ is almost the same at low temperature regime in FM and AFM. This is due to almost the same magnitude of Zeeman gap, as shown in Figs. 5(a) and 5(b) for the FM and AFM cases, respectively, which is further confirmed by the linear spin-wave theory in Sec. III.3. Figure 5(c) also shows $\tilde{J}_{S}$ for the FM Heisenberg model at $B=0.5$ for comparison; the magnitude is much smaller than the result for $B=0.01$ shown in Fig. 4(c) because of the larger Zeeman gap, but it is comparable to those for the FFM states in the FM and AFM Kitaev models.
@@ III.2 Perturbation analysis
To examine the origin of the different sign of $\tilde{J}_{S}$ between the FM and AFM KSLs in Fig. 4(c), we calculate the tunnel spin current in terms of Majorana fermions by using the perturbation theory with respect to the magnetic field. Following the seminal paper by Kitaev [24], we represent each spin $S=1/2$ operator by four Majorana fermion operators as $S^{\gamma}_{j}=ib^{\gamma}_{j}c_{j}/2$, where $b^{\gamma}_{j}$ and $c_{j}$ are the Majorana fermion operators. By the transformation, the original spin degree of freedom is fractionalized into two types of quasiparticles: itinerant Majorana fermions and localized $Z_{2}$ fluxes. The latter is defined by a product of the Majorana fermion operators $b_{j}^{\gamma}$, being local conserved quantities that take on values of ±$\pm 1$. The sector with all $+1$ fluxes contains the ground state of the system and is called flux free sector. In the spinless Majorana fermion representation, the Kitaev Hamiltonian at zero field in the flux free sector is represented as
@@
with the spinless Majorana fermion operators $c_{j}$. Based on the general framework of the perturbation theory [94], we obtain the effective operator of $S^{+}_{j_{0}}$ in the flux free sector through a transformation by $\Gamma:=PP_{0}(P_{0}PP_{0})^{-1/2}$, where $P_{0}$ and $P$ are projectors to the flux free sector of the unperturbed Hamiltonian [Eq. (1) for $B=0$] and to the states of the full Hamiltonian that are adiabatically connected to those in the flux free sector at $B=0$, respectively. By expanding up to the second order with respect to $B=|{\bf B}|$, the effective operator is obtained as
@@
$\left.X_{+}(\omega)\right|_{S^{\pm}_{j_{0}}\to s^{\pm}_{j_{0}},\mathcal{H}\to\mathcal{H}_{0}}=\delta\sum_{\ell}\frac{|\langle\ell|s^{-}_{j_{0}}|\emptyset\rangle|^{2}-|\langle\ell|s^{+}_{j_{0}}|\emptyset\rangle|^{2}}{(E_{\ell}-E_{\emptyset}-\omega)^{2}+\delta^{2}}$ (13)
@@
$\tilde{J}_{S}\sim\frac{B^{3}}{|K|^{3}}\tilde{j}_{S}$ (14)
@@
Figure 7: Results of the perturbation analysis for the FM and AFM Kitaev models in the weak magnetic field limit with ${\bf B}\parallel a$, ${\bf B}\parallel b$, and ${\bf B}\parallel c$: (a) $\omega$ dependences of $x_{+}(\omega)$ [Eq. (16)] and (b) temperature dependences of $\tilde{j}_{S}$ [Eq. (15)]. The lines represent the full results including all the contributions from $s^{+(1)}$ and $s_{1,2}^{+(2)}$, and the hatched regions represent the results ignoring $s^{\pm(2)}_{2}$. In the latter hatched regions, ${\bf B}\parallel a$ and ${\bf B}\parallel b$ give the same results.
@@
The perturbation analysis provides a useful insight on the origin of the sign difference of the tunnel spin current between the FM and AFM cases. When we neglect the four Majorana contributions of $s_{2}^{\pm(2)}$, one can show that $x_{+}(\omega)$ in the FM and AFM cases is identical in the absolute value, but the signs are opposite. This is because the $\mathcal{O}(B^{3})$ contributions in Eq. (13) from the cross terms between $s^{\pm(1)}$ and $s^{\mp(2)}_{1}$ change their sign by $K\to-K$, i.e., $\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle\to-\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle$, due to the sign change in the eigenstates of Majorana fermions (see Appendix B). Consequently, the interchange of FM and AFM $K$ causes a sign reversal of $\tilde{j}_{S}$ in Eq. (15). The contributions with neglecting $s_{2}^{\pm(2)}$ are plotted by the hatched regions in Fig. 7. The result indicates that the contributions from the quadratic terms of the Majorana operators, $s^{\pm(1)}$ and $s^{\pm(2)}_{1}$, are dominant in $\tilde{j}_{S}$, and moreover, are the main cause of the sign switching of the tunnel spin current between the FM and AFM cases. The asymmetry between the two cases is due to the quartic terms in $s_{2}^{\pm(2)}$, leading to the faster growth of the tunnel spin current with temperature in AFM than FM, which is also consistent with the TDVP results in Fig. 4(c).
Let us make a remark on the $B$ dependence of the tunnel spin current. The perturbation theory in the low-field limit predicts $\tilde{j}_{S}\propto B^{3}$, but $\tilde{J}_{S}$ obtained by TDVP appears not to follow this asymptotic form; we find that $\tilde{J}_{S}$ at $B=B_{\rm KSL}/2$ is about half of that at $B=B_{\rm KSL}$ (not shown), suggesting that $\tilde{J}_{S}$ is roughly proportional to $B$. There are at least two possible reasons for this discrepancy. One is the contributions from excited states beyond the flux-free sector, which are neglected in the current perturbation analysis. The intermediate states [$|\ell\rangle$ in Eq. (13)] including flux excitations may give $\mathcal{O}(B)$ contributions, although they will be suppressed exponentially in the low-field limit because of the finite flux gap. If this is the case, the perturbation theory predict $\tilde{j}_{S}\simeq pB^{3}+qBe^{-\Delta_{\rm f}/T}$, where $\Delta_{\rm f}$ is the flux gap $\simeq 0.06|K|$ [24]. The other possibility is the limited system sizes in the TDVP simulations. Larger size simulations allow us to distinguish the $B$ dependence by accessing lower-field regions with higher precision by taking smaller damping factor $\delta$ in Eq. (3). While such calculations require huge computational costs and are left for future study, this study comprehensively suggests that the coefficient $p$ and $q$ are both positive (negative) for the FM (AFM) KSL.
@@ III.3 Linear spin-wave theory
Figure 8: Results by the linear spin-wave theory for the (a,b) FM and (c,d) AFM Kitaev models at $B=B_{\rm FFM}$ with four field directions ${\bf B}\parallel a$, ${\bf B}\parallel b$, ${\bf B}\parallel c$, and ${\bf B}\parallel z$. (a) and (c) show the $\omega$ dependences of $X_{+}(\omega)$ for the FM and AFM cases, respectively. (b) and (d) show the corresponding temperature dependences of the tunnel spin current $\tilde{J}_{S}$.
@@
Figure 8 shows the results for $B=B_{\rm FFM}$ (see Fig. 3) with four different field directions. We find that $X_{+}(\omega)$ is always positive for both FM and AFM cases, as shown in Figs. 8(a) and 8(c), respectively. These behaviors are in qualitative agreement with the TDVP simulation results shown in Figs. 5(a) and 5(b). The positive $X_{+}(\omega)$ leads to the tunnel spin current with positive sign as shown in Figs. 8(b) and 8(d). This is also in agreement with the TDVP simulation results in Fig. 5(c). We note, however, that the gap in the spectrum of $X_{+}(\omega)$ is independent of the field directions in the linear spin-wave theory, while it is considerably different between ${\bf B}\parallel z$ and ${\bf B}\parallel c$ in the TDVP simulation.
@@ IV.1 Sign of spin current in Kitaev spin liquids
The real-time dynamics simulations based on TDVP reveal the sign switching of the tunnel spin current between the FM and AFM KSLs (Sec. III.1). This behavior is successfully reproduced by the perturbation theory in the Majorana representation (Sec. III.2). These findings indicate that the spin current in the KSL is predominantly carried by the itinerant Majorana fermions. More importantly, they suggest that the itinerant Majorana fermions contribute to the spin current with (up-)down-spin-like nature in the (AFM) FM KSL where the sign of the tunnel spin current is (negative) positive, even though they do not have spin angular momentum.
The negative SSE that we find for the AFM KSL appears to be common to that for a quantum spin liquid in a quasi-one-dimensional antiferromagnet [62]. This implies common physics in the spin current carried by fractional excitations in AFM quantum spin liquids in one and two dimensions, even though the fractional quasiparticles have different nature: The spin angular momentum carried by a spinon in the one-dimensional quantum spin liquid is half-integer quantized $\Delta S=1/2$, while that by a Majorana fermion in the KSL is not. Our results, however, indicate that the FM quantum spin liquid, which is peculiar to the Kitaev model, can induce a tunnel spin current with positive sign, suggesting that the SSE is useful to identify the nature of fractional quasiparticles in quantum spin liquids. We emphasize that our study reveals, for the first time to our knowledge, that the SSE in quantum spin liquids can be either positive or negative by analyzing both FM and AFM Kitaev models in two dimensions. Toward deeper understanding of the SSE in quantum spin liquids, further comparative studies are necessary for different types of quantum spin liquids in different dimensions. The simulation method for the real-time dynamics used in this study provides a powerful method to study generic models and awaits wider applications in the future.
Meanwhile, the SSE for the FM KSL has the same positive sign as that carried by conventional magnons in the FFM state. Thus, a way to distinguish between the two is needed. We will discuss this point in the next subsection.
@@ IV.2 Magnetic field dependence of tunnel spin current
Figure 9: Schematics of the magnetic field dependence of the tunnel spin current caused by the SSE at low temperature for (a) the FM and AFM Kitaev models and (b) the FM Heisenberg model.
@@
Based on the results by the real-time dynamics simulations, the perturbation theory, and the linear spin-wave theory, we show a schematic of the magnetic field dependence of the tunnel spin current caused by the SSE for the Kitaev model in Fig. 9(a). On one hand, in the low-field KSL region, the tunnel spin current increases in amplitude with increasing the field strength, with a different sign between the FM and AFM cases. As discussed in the end of Sec. III.2, a more detailed analysis is needed to fully elucidate the behavior in the low-field limit. On the other hand, in the high-field FFM regime, the tunnel spin current is positive for both FM and AFM cases and decreases with increasing the field strength as shown in Fig. 9(a). This is similar to the FM Heisenberg model [Fig. 9(b)]. Thus, these results imply that the tunnel spin current is maximized between the KSL and FFM states in the FM case, while it should change its sign in the AFM cases, as schematically shown in Fig. 9(a). However, further studies are needed to clarify the field dependence in those transient regions, including the another intermediate SL region in the AFM case (see Fig. 3).
@@
Figure 10: Field-angle dependences of the tunnel spin current when the magnetic field is rotated in the $a$-$c$ plane obtained by (a) the perturbation analysis for low $B$ and (b) the linear spin-wave theory at $B=B_{\rm FFM}$. The angle $\theta$ is measured from the $a$ axis, and the data are calculated at $T=0.05$.